# Evaluating Evaluation Metrics – The Mirage of Hallucination Detection

### v1

@@ A.2 Models
  • OPT Zhang et al. (2022) – 125M, 350M, 1.3B, 2.7B, 6.7B, 13B, 30B, 66B
  • Llama 2Touvron et al. (2023) – 7B, 13B, 70B, 7B-chat, 13B-chat, and 70b-chat
  • Llama 3 and 3.1 Dubey et al. (2024) – 3B, 70B, 3B-Instruct, 70B-Instruct
  • Phi Gunasekar et al. (2023); Li et al. (2023b); Abdin et al. (2024) – Phi-3-small-8k-instruct, Phi-3-mini-4k-instruct, Phi-3-medium-4k-instruct
  • Gemma Gemma (2024) – gemma-2b, gemma-7b, gemma-2b-it, gemma-7b-it, gemma-1.1-2b-it, gemma-1.1-7b-it
  • OLMo Groeneveld et al. (2024) – 1B, 7B, 1B-Instruct, and 7B-Instruct
@@ A.3 Prompts
[…]
@@ Appendix C Extended Discussions
C.1 Most Metrics Exhibit Poor Inter-Metric Correlation and Misalignment with Human Judgment
@@
C.2 Mode-Seeking Decoding Hallucinate less than Sampling-based Approaches
The box plots in Figures 7, 8, and 9 illustrate the performance of various decoding techniques across different metrics. The decoding methods considered include greedy, beam search ($b=3$), ancestral, top-k ($k=40$), and top-p ($p=0.95$). Models are grouped by parameter size into the following bins: >0.5,>4,>20,>70{>0.5,>4,>20,>70}> 0.5 , > 4 , > 20 , > 70 billion parameters. Overall, greedy and beam search consistently outperform sampling-based methods. However, this trend breaks for BertScore and K-BertScore in the case of FaithDial. We hypothesize that this is possibly due to the model’s limited capacity, which may lead to repetitive or degenerate outputs, as observed in previous studies Holtzman et al. (2020). Other metrics such as Knowledge-F1, $\mathcal{Q}^{2}$ token F1, MSP, and Perplexity adhere to the trend.
The heatmaps in Figures 10 and 11 show the p-values for pairwise significance tests between the decoding methods. Except for the consistency score, greedy and beam search consistently outperform sampling-based methods with statistically significant results. These findings further confirm that probability-maximization decoding methods help reduce hallucinations, particularly in knowledge-grounded tasks.
@@ C.2 Mode-Seeking Decoding Hallucinate less than Sampling-based Approaches
C.3 Parameter Scaling does not Necessarily Improve Hallucination Metrics
Figures 12 and 13 illustrate the performance of NLG, token-overlap, and uncertainty-based metrics as language model parameters scale. While Rouge-L, $\mathcal{Q}^{2}$ token F1, MSP, and perplexity all improve with model size, other metrics do not show a consistent pattern. Figure 14 presents the p-values for pairwise significance tests across different model sizes, revealing that BertScore shows no improvement as the model size increases.

### v2

@@ A.2 Models
  • OPT Zhang et al. (2022): 125M, 350M, 1.3B, 2.7B, 6.7B, 13B, 30B, 66B
  • Llama 2 Touvron et al. (2023): 7B, 13B, 70B, 7B-chat, 13B-chat, and 70b-chat
  • Llama 3 and 3.1 Dubey et al. (2024): 3B, 70B, 3B-Instruct, 70B-Instruct
  • Phi Gunasekar et al. (2023); Li et al. (2023b); Abdin et al. (2024): Phi-3-small-8k-instruct, Phi-3-mini-4k-instruct, Phi-3-medium-4k-instruct
  • Gemma Gemma (2024): gemma-2b, gemma-7b, gemma-2b-it, gemma-7b-it, gemma-1.1-2b-it, gemma-1.1-7b-it
  • OLMo Groeneveld et al. (2024): 1B, 7B, 1B-Instruct, and 7B-Instruct
@@ A.3 Prompts
The GPT-4 evaluator prompts for Begin, TruthfulQA, and FaithDial are outlined in Tables 3, 4, and 5, respectively. We use the OpenAI’s gpt-4o-mini model. The evaluator selects from the options $\{\texttt{A, B, C, D, E}\}$, with options B and D identified as hallucinated responses, while the others are deemed benign. These prompts are based on the factuality template from the Autoevals library [^2] https://github.com/braintrustdata/autoevals/. For HaluEval, we leverage the pre-existing templates provided by Li et al. (2023a).
@@ Appendix C Extended Discussions
C.1 Most Metrics Exhibit Poor Alignment with Human Judgment
Dataset ROUGE-L SacreBleu BertScore Knowledge-F1 $\mathcal{Q}^{2}$ token-F1
Begin CMU - - - $0.72$ $0.70$
Begin TC - - - $0.75$ $0.74$
Begin WoW - - - $0.43$ $0.53$
HaluEval Dial $0.31$ $0.32$ $0.31$ $0.59$ $0.53$
HaluEval QA $0.30$ $0.54$ $0.31$ $0.83$ $0.81$
Table 9: PRAUC scores between rest of the hallucination metrics and human annotations.
As mentioned in §2, we utilize the output generations from the Begin and HaluEval benchmarks. Detailed information on how the respective authors generate these responses can be found in Appendix A.1. Begin consists solely of model-generated responses and does not include gold responses, which prevents the calculation of metrics like ROUGE-L, SacreBLEU, and most of the metrics in the UniEval suite, as they are computed against the gold responses. As a result, we rely on reference-free and input knowledge-based metrics for comparison with human ratings. Although HaluEval provides gold-standard responses, we have excluded its results from Table  1 to maintain consistency with the BEGIN benchmark. Table 9 provides the results (PRAUC scores) for the remaining metrics. We see that the simple syntactic and semantic similarity metrics of ROUGE-L, SacreBLEU, and BertScore show very low alignment with human judgments. Knowledge-F1 and $\mathcal{Q}^{2}$ token-F1 yeild similar scores to Knowledge-BertScore and $\mathcal{Q}^{2}$-NLI F1 score.
@@ C.1 Most Metrics Exhibit Poor Alignment with Human Judgment
C.2 Why is the Inter-Metric Correlation Weak?
Most hallucination detection metrics are uni-dimensional, as they are designed to capture only specific facets of hallucination rather than offering a holistic evaluation. This design limitation leads to low inter-metric correlation, as different metrics often emphasize fundamentally different properties of hallucinated content. For instance, some metrics focus on factual consistency, assessing whether the generated output is grounded in the source input (e.g., question, context, or prompt). Others may concentrate on fluency, semantic similarity, or entity-level accuracy. Because these properties are orthogonal, a model might score well on one metric while performing poorly on another.
GPT-4 based evaluation is a better metric for detecting hallucinations because unlike automated metrics that rely on predefined heuristics (e.g., n-gram overlap, embeddings, or NLI classifiers), GPT-4 can assess nuanced errors, infer missing knowledge, and detect inconsistencies in a way that aligns closely with human judgment, as it considers various factors such as coherence, commonsense reasoning, and factual grounding, to name a few (OpenAI, 2024). Here is why it shows a weak correlation with different metrics:
  • GPT-4 vs. N-gram Overlap (Rouge-L, SacreBLEU, and Knowledge-F1): GPT-4 assesses meaning and factuality beyond simple word overlap, whereas these metrics only measure surface-level similarity. A hallucinated response can have a high n-gram overlap with a reference while still being incorrect, leading to false positives. Conversely, correct but reworded responses can be penalized, leading to false negatives. GPT-4’s reasoning capabilities makes it more flexible than rigid n-gram matching.
  • GPT-4 vs. Semantic Similarity (BERTScore and K-BERTScore): These metrics measure embedding similarity but do not verify factual accuracy. Two sentences can be semantically close while differing in factual correctness. GPT-4 can assess fine-grained factual inconsistencies that semantic similarity models miss, such as incorrect numerical values or subtly misleading statements.
  • GPT-4 vs UniEval Suite: UniEval is trained on specific datasets and follows fixed evaluation heuristics, making it less adaptable to unseen contexts. GPT-4 dynamically evaluates responses using broad-world knowledge and deeper reasoning, leading to higher accuracy in detecting nuanced hallucinations.
  • GPT-4 vs. $\mathcal{Q}^{2}$: It relies on question generation and answer extraction, which introduces cascading errors if the generated questions are poorly framed or if the extraction mechanism fails. Moreover, it may overlook implicit hallucinations that do not map neatly to question-answer pairs, whereas GPT-4 can reason about implicit information.
  • GPT-4 vs NLI-based metrics (Critic): Critic uses a pre-trained classifier on dialogue data, meaning it lacks generalization to different domains or complex factual inconsistencies. NLI models often misinterpret negations, indirect claims, and paraphrased statements, leading to misclassifications that GPT-4 would avoid.
C.3 Mode-Seeking Decoding Hallucinate less than Sampling-based Approaches
The box plots in Figures 8, 9, and 10 illustrate the performance of various decoding techniques across different metrics. The decoding methods considered include greedy, beam search ($b=3$), ancestral, top-k ($k=40$), and top-p ($p=0.95$). Models are grouped by parameter size into the following bins: >0.5,>4,>20,>70{>0.5,>4,>20,>70} billion parameters. Overall, greedy and beam search consistently outperform sampling-based methods. However, this trend breaks for BertScore and K-BertScore in the case of FaithDial. We hypothesize that this is possibly due to the model’s limited capacity, which may lead to repetitive or degenerate outputs, as observed in previous studies Holtzman et al. (2020). Other metrics such as Knowledge-F1, $\mathcal{Q}^{2}$ token F1, MSP, and Perplexity adhere to the trend.
The heatmaps in Figures 11 and 12 show the p-values for pairwise significance tests between the decoding methods. Except for the consistency score, greedy and beam search consistently outperform sampling-based methods with statistically significant results. These findings further confirm that probability-maximization decoding methods help reduce hallucinations, particularly in knowledge-grounded tasks.
@@ C.3 Mode-Seeking Decoding Hallucinate less than Sampling-based Approaches
C.4 Parameter Scaling does not Necessarily Improve Hallucination Metrics
Figures 13 and 14 illustrate the performance of NLG, token-overlap, and uncertainty-based metrics as language model parameters scale. While Rouge-L, $\mathcal{Q}^{2}$ token F1, MSP, and perplexity all improve with model size, other metrics do not show a consistent pattern. Figure 15 presents the p-values for pairwise significance tests across different model sizes, revealing that BertScore shows no improvement as the model size increases.
@@ C.4 Parameter Scaling does not Necessarily Improve Hallucination Metrics
In summary, to the best of our knowledge, our work is the first to comprehensively evaluate a wide range of hallucination detection metrics at scale, across multiple datasets, model families, model sizes, decoding strategies, and training methods. While Finding 3 have been established in prior studies, such as Dziri et al. (2021) and Ouyang et al. (2022), they lack the robustness provided by our analysis, as they were not tested across the diverse dimensions that we explore. Our work offers a more thorough and holistic assessment, demonstrating that these findings indeed hold true across different settings and providing deeper insights for ML and NLP practitioners about which metrics perform best under various conditions. Moreover, to the best of our knowledge, none of the previous works have concretely shown the emergence of finding 4. Lastly, while some of these findings might seem obvious at first, we believe scientific research is often exactly around such contributions - transforming intuitive observations into a robust, evidence-backed understanding, advancing the field with concrete, reproducible findings.